# Transforming Policy-Car Swerving for Mitigating Stop-and-Go Traffic Waves: A Practice-Oriented Jam-Absorption Driving Strategy

### v1

@@ IV Theoretical analysis of the JAD
$J_{\mathrm{dis}}$ and $J_{\mathrm{dur}}$ are linearly related through distance–time–speed relationship, and are determined by five parameters, i.e., $v^{*}$, $v_{\mathrm{t}}$, $\Delta_{\mathrm{w}}$, $w$, $v_{\mathrm{w}}$. Therefore, analyzing the impact of system parameters on $J_{\mathrm{dis}}$ is sufficient to characterize the behavior of the JAD process, without loss of generality.
To facilitate the interpretation of the theoretical results, we summarize empirically plausible parameter ranges related to traffic jam waves and JAD vehicles in Table I. These values are consistent with real-world traffic observations and serve as practical references for the subsequent analysis of the impact of various factors on JAD implementation.
@@
JAD speed $v^{*}$ 40 80 km/h Desired JAD speed that may not trigger secondary waves, considering traffic flow stability
Inflow traffic speed $v_{\mathrm{t}}$ 80 120 km/h Speed of traffic approaching the stop-and-go traffic wave from upstream
Wave temporal width $\Delta_{\mathrm{w}}$ 20 100 s Temporal width of the stop-and-go traffic wave
Wave speed $w$ 20-20 10-10 km/h Wave speed of the stop-and-go traffic wave propagating upstream
In-wave speed $v_{\mathrm{w}}$ 0 20 km/h Traffic speed inside the moving queue associated with the stop-and-go traffic wave
IV-A The impact of JAD speed
@@
$\frac{\partial J_{\mathrm{dis}}}{\partial v^{*}}=\frac{w\,\Delta_{\mathrm{w}}\,(v_{\mathrm{t}}-v_{\mathrm{w}})\,v_{\mathrm{t}}}{(w-v_{\mathrm{w}})\,(v_{\mathrm{t}}-v^{*})^{2}}$ (22)
@@
A rough estimate using typical parameter ranges from Table I shows that $\partial J_{\mathrm{dis}}/\partial v^{*}$ is on the order of $10^{1}\sim 10^{2}~\mathrm{m/(km/h)}$, which corresponds to a JAD duration on the order of $10^{2}\sim 10^{3}~\mathrm{s}$. This implies that an increase of 1 km/h in the JAD speed may require an additional $\sim10$0 m of execution distance. The results of the numerical simulations shown in Figure 6 indicate that $J_{\text{dis}}$ ranges from approximately 2 km to more than 12 km, given moderate values of other parameters.
@@
IV-B The impact of Inflow Traffic Speed
@@
$\frac{\partial J_{\mathrm{dis}}}{\partial v_{\mathrm{t}}}=\frac{v^{*}w\Delta_{\mathrm{w}}(w-v^{*})(v^{*}-v_{\mathrm{w}})}{(v^{*}-w)(w-v_{\mathrm{w}})(v_{t}-v^{*})^{2}}$ (23)
@@
Physically, as illustrated in Figure 7, a faster inflow (dashed line) results in a shorter distance and a shorter duration of the JAD process, under the constraint that all other variables are held constant. The nonlinear dependence of $J_{\mathrm{dis}}$ on $v_{\mathrm{t}}$ is dominated by the inverse-square term $(v_{\mathrm{t}}-v^{*})^{-2}$ in (23), implying the marginal reduction diminishes as $v_{\mathrm{t}}$ increases.
A rough order-of-magnitude estimate based on typical parameter ranges in Table I indicates that $\partial J_{\mathrm{dis}}/\partial v_{\mathrm{t}}$ is on the order of $10^{1}\sim-10^{2}~\mathrm{m/(km/h)}$, which corresponds to $J_{\mathrm{dur}}$ on the order of $10^{0}\sim 10^{1}~\mathrm{min}$. The results of the numerical simulations shown in Figure 6 indicate that $J_{\text{dis}}$ ranges from approximately 4 km to 7.5 km, given moderate values of other parameters.
@@
IV-C The Impact of Wave Temporal Width
@@
$\frac{\partial J_{\mathrm{dis}}}{\partial\Delta_{\mathrm{w}}}=\frac{v^{*}w(v_{\mathrm{t}}-v_{\mathrm{w}})}{(w-v_{\mathrm{w}})(v_{\mathrm{t}}-v^{*})}$ (24)
@@
Increasing the wave duration from its lower to upper bound in Table I typically increases $J_{\mathrm{dis}}$ by an order of $10^{1}\sim 10^{2}~\mathrm{m}$, corresponding to $J_{\mathrm{dur}}$ on the order of $10^{0}\sim 10^{1}~\mathrm{min}$. The simulation results in Figure 8 further indicate that, for moderate values of the other parameters, $J_{\mathrm{dis}}$ ranges from approximately $2~\mathrm{km}$ to $8~\mathrm{km}$.
@@
IV-D The Impact of Wave Speed
@@
Varying the wave speed within the range reported in Table I typically leads to changes in $J_{\mathrm{dis}}$ on the order of $10^{2}\sim 10^{3}~\mathrm{m}$, and in $J_{\mathrm{dur}}$ on the order of $10^{1}\sim 10^{2}~\mathrm{s}$. The simulation results in Figure 9 further show that, for moderate values of the other parameters, the range of $J_{\mathrm{dis}}$ is between 4 km and 5.4 km, and the range of $J_{\mathrm{dur}}$ is between 65 s and 90 s. Compared with the ranges resulted by other parameters, the wave speed’s impact on the JAD process is relatively small.
@@
IV-E The Impact of In-Wave Traffic Speed
@@
Physically, when the vehicles inside the stop-and-go wave move faster, the JAD Triangle becomes smaller, and the JAD process is shortened accordingly. This means that, compared with a stop-and-go wave, a slow-moving-and-go wave is easier to mitigate.
@@ IV-F Summary of Parameter Sensitivity
Based on the preceding theoretical analysis, the sensitivity of the JAD implementation to key system parameters can be summarized as follows and in Table II.
  • The impacts of JAD speed $v^{*}$ and wave width $\Delta_{\mathrm{w}}$ on $J_{\mathrm{dis}}$ and $J_{\mathrm{dur}}$ are positive. That is, higher $v^{*}$ and larger $\Delta_{\mathrm{w}}$ lead to a longer JAD process, while fixing all other paremeters.
  • The impacts of inflow traffic speed $v_{t}$, wave speed $w$, and in-jam traffic speed $v_{\mathrm{w}}$ on $J_{\mathrm{dis}}$ and $J_{\mathrm{dur}}$ are negative. It indicates that a higher $v_{t}$, a more slowly propagating wave (smaller $|w|$), and higher $v_{\mathrm{w}}$ require a shorter JAD process.
@@
  • Considering the moderate values of the other parameters in Table I, the relative influence of each parameter on $J_{\mathrm{dis}}$ and $J_{\mathrm{dur}}$, from strongest to weakest, is:
@@
1 JAD speed $v^{*}$ ++ [2, 13] [50, 160]
2 Wave width $\Delta_{\mathrm{w}}$ ++ [1, 8] [30, 140]
3 Wave speed $w$ - [3, 9] [50, 150]
4 Inflow speed $v_{\mathrm{t}}$ - [4, 7.5] [60, 130]
5 In-wave speed $v_{\mathrm{w}}$ - [4, 5.4] [60, 90]
@@ V-A Simulation Scenario
We employ SUMO to simulate traffic on a single-lane straight freeway. The road length is 8 km, with a speed limit of 90 km/h. The simulation runs for 1600 s with a time step of 1 s. Vehicles enter the network at the highway entrance every 1.5 s [^4] Such a persistently high demand is effective in ensuring the continuous growth of a stop-and-go wave. We also experimented with more stochastic demand patterns, but failed to produce a self-sustaining wave. with an initial speed of 54 km/h, and they soon accelerate to their maximum speed. The Krauss car-following model embedded in SUMO is employed, and the main parameters are summarized in Table III. All the code has been open-sourced on our GitHub repository (https://github.com/gotrafficgo).
@@
Figure 11: Simulatied self-sustaining stop-and-go wave.
@@ V-B1 JAD Speed $v^{*}$
We reduce the speed of the first vehicle to 30, 40, 50, and 60 km/h, respectively, and simulate each scenario to observe which speeds do not trigger secondary waves. As shown in Figure 12, speeds of 30 and 40 km/h result in traffic breakdown, and are therefore unsuitable as JAD speeds. In contrast, speeds of 50 km/h or higher, such as 50 and 60 km/h, can serve as reliable JAD speeds. However, as discussed in Section IV-A, increasing the JAD speed also increases the time and space required to execute the JAD task, implying a longer operational period.
@@
V-B2 Wave temporal width $\Delta_{\mathrm{w}}$
@@
Typically, a stationary detector monitors the traffic and measures the speed of passing vehicles in real time. Therefore, we assume that a stop-and-go wave has formed when the observed speed falls below a certain threshold, and continues to remain low for a period of time. This approach is feasible because the target of this study is a standalone stop-and-go wave. In contrast, for periodic stop-and-go waves generated by fixed bottlenecks, this method may not be suitable, and the proposed JAD strategy is not designed for such traffic waves.
In this scenario, we set the threshold to 36 km/h and the holding period to 30 s. As shown in Figure 13, the resulting time-space points $F$ and $E$ effectively capture the stop-and-go wave. It is also noted that, due to the asymmetric behavior of vehicles traversing a stop-and-go wave [4, 20, 27], the recovery process from a slowly moving queue is longer than the deceleration process. As will be demonstrated later in Section V-C with simulation results, a slight overestimation of the wave speed, while sacrificing some efficiency, not only does not compromise the goal of halting the propagation of the stop-and-go wave, but may also improve the robustness of the JAD strategy.
@@
V-B3 In-wave Speed $\Delta_{\mathrm{w}}$
As shown in Figure 13, from the practical perspective of a stationary roadside detector, the speed variation through a stop-and-go wave exhibits a gradual decrease followed by a gradual increase. As a result, the stationary roadside detector may be unable to accurately capture the in-wave speed.
Here, we take the average speed between $t_{F}$ and $t_{E}$ as the in-wave speed $\Delta_{\mathrm{w}}$. If the minimum value were used instead, it could be underestimated.
V-B4 Wave Speed $w$
@@
Here, from the simulation, we set the wave speed to w=15w=-15 km/h.
@@ V-B7 Opportunity of Dispatching a JAD Vehicle
In the police-car-swerving scenario, the police officer is expected to identify an appropriate gap for merging. If merging from an on-ramp, the overall process would be as natural as typical daily highway merging behavior. With the siren activated, the required gap may be even smaller than that under normal conditions.
@@ V-C Simulation Results
The results could be further improved by fine-tuning the parameters. For example, a more accurate estimation of the wave width could enhance efficiency by enabling a better prediction of the ending point $C$ of the JAD strategy. However, this refinement was not implemented because, in practice, traffic conditions cannot be perfectly replicated, and obtaining an exact value is challenging.
Note that it is not necessary to compare travel time or emissions with the base case in Figure 11, since once the stop-and-go wave is suppressed, the following vehicles naturally experience shorter travel times and lower emissions when operating under smoother traffic conditions without falling into a stop-and-go process.
@@
Figure 16 presents two typical cases in which the JAD vehicle fails to suppress the stop-and-go wave, as a secondary wave is triggered. In Figure 16(a), the JAD speed is set to $v^{*}$ = 35 km/h. As indicated by the stability test in Figure 12, such a low speed is insufficient to prevent traffic breakdown. In Figure 16(b), the JAD vehicle fails to avoid being captured by the stop-and-go wave due to an underestimation of the wave width. As a result, the stop-and-go wave penetrates the JAD trajectory and continues to propagate downstream.

### v2

@@ IV Theoretical analysis of the JAD
$J_{\mathrm{dis}}$ and $J_{\mathrm{dur}}$ are linearly related through distance-time-speed relationship, and are determined by five parameters, i.e., $v^{*}$, $v_{\mathrm{t}}$, $\Delta_{\mathrm{w}}$, $w$, $v_{\mathrm{w}}$. Therefore, analyzing the impact of the parameters on $J_{\mathrm{dis}}$ is sufficient to characterize the behavior of the JAD process, without loss of generality.
To facilitate the interpretation of the theoretical results, we summarize empirically plausible parameter ranges related to stop-and-go waves and JAD vehicles in Table I. These values are consistent with real-world traffic observations and serve as practical references for the subsequent analysis of the impact of various factors on JAD implementation.
@@
JAD speed ($v^{*}$) 40 80 km/h Desired JAD speed that may not trigger secondary waves, considering traffic flow stability
Inflow traffic speed ($v_{\mathrm{t}}$) 80 120 km/h Speed of traffic approaching the stop-and-go traffic wave from upstream
Wave temporal width ($\Delta_{\mathrm{w}}$) 20 100 s Temporal width of the stop-and-go traffic wave
Wave speed ($w$) 20-20 10-10 km/h Wave speed of the stop-and-go traffic wave propagating upstream
In-wave speed ($v_{\mathrm{w}}$) 0 20 km/h Traffic speed inside the moving queue associated with the stop-and-go traffic wave
IV-A Impact of JAD Speed
@@
$\frac{\partial J_{\mathrm{dis}}}{\partial v^{*}}=\frac{w\,\Delta_{\mathrm{w}}\,(v_{\mathrm{t}}-v_{\mathrm{w}})\,v_{\mathrm{t}}}{(w-v_{\mathrm{w}})\,(v_{\mathrm{t}}-v^{*})^{2}}.$ (22)
@@
A rough estimate using typical parameter ranges from Table I shows that $\partial J_{\mathrm{dis}}/\partial v^{*}$ is on the order of $10^{1}\sim 10^{2}$ m, which corresponds to a JAD duration on the order of $10^{2}\sim 10^{3}~\mathrm{s}$. This implies that an increase of 1 km/h in the JAD speed may require an additional $\sim10$0 m of execution distance. The results of the numerical simulations shown in Figure 5 indicate that $J_{\text{dis}}$ ranges from approximately 2 km to more than 12 km and $J_{\text{dur}}$ ranges from approximately 50 s to more than 160 s, given moderate values of other parameters.
@@
IV-B Impact of Inflow Traffic Speed
@@
$\frac{\partial J_{\mathrm{dis}}}{\partial v_{\mathrm{t}}}=\frac{v^{*}w\Delta_{\mathrm{w}}(w-v^{*})(v^{*}-v_{\mathrm{w}})}{(v^{*}-w)(w-v_{\mathrm{w}})(v_{t}-v^{*})^{2}}.$ (23)
@@
Physically, as illustrated in Figure 6, a faster inflow (red dashed line) results in a shorter distance and a shorter duration of the JAD process, under the constraint that all other variables are held constant. The nonlinear dependence of $J_{\mathrm{dis}}$ on $v_{\mathrm{t}}$ is dominated by the inverse-square term $(v_{\mathrm{t}}-v^{*})^{-2}$ in (23), implying the marginal reduction diminishes as $v_{\mathrm{t}}$ increases.
A rough order-of-magnitude estimate based on typical parameter ranges in Table I indicates that $\partial J_{\mathrm{dis}}/\partial v_{\mathrm{t}}$ is on the order of $10^{1}\sim 10^{2}$ m, which corresponds to $J_{\mathrm{dur}}$ on the order of $10^{1}\sim 10^{2}s$ s. The results of the numerical simulations shown in Figure 5 indicate that $J_{\text{dis}}$ ranges from approximately 4 to 7.5 km and $J_{\text{dur}}$ ranges from approximately 60 to 130 s, given moderate values of other parameters.
@@
IV-C Impact of Wave Temporal Width
@@
$\frac{\partial J_{\mathrm{dis}}}{\partial\Delta_{\mathrm{w}}}=\frac{v^{*}w(v_{\mathrm{t}}-v_{\mathrm{w}})}{(w-v_{\mathrm{w}})(v_{\mathrm{t}}-v^{*})}.$ (24)
@@
Increasing the wave duration from its lower to upper bound in Table I typically increases $J_{\mathrm{dis}}$ by an order of $10^{1}\sim 10^{2}$ m, corresponding to $J_{\mathrm{dur}}$ on the order of $10^{1}\sim 10^{2}$ s. The simulation results in Figure 7 further indicate that, for moderate values of the other parameters, $J_{\mathrm{dis}}$ ranges from approximately 2 to 8 km, and $J_{\mathrm{dur}}$ ranges from approximately 30 to 140 s.
@@
IV-D Impact of Wave Speed
@@
Varying the wave speed within the range reported in Table I typically leads to changes in $J_{\mathrm{dis}}$ on the order of $10^{2}\sim 10^{3}$ m, and in $J_{\mathrm{dur}}$ on the order of $10^{1}\sim 10^{2}$ s. The simulation results in Figure 8 further show that, for moderate values of the other parameters, the range of $J_{\mathrm{dis}}$ is between 4 and 5.4 km, and the range of $J_{\mathrm{dur}}$ is between 65 and 90 s. Compared with the ranges resulted by other parameters, the wave speed’s impact on the JAD process is relatively small.
@@
IV-E Impact of In-Wave Traffic Speed
@@
Physically, when the vehicles inside the stop-and-go wave move faster, the JAD Triangle becomes smaller, and the JAD process is shortened accordingly. This means that, compared with a stop-and-go wave, a slow-moving-and-go wave is easier to mitigate, which is consistent with the common sense.
@@ IV-F Summary of Parameter Sensitivity
Based on the preceding theoretical analysis, the sensitivity of the JAD implementation to key parameters can be summarized as follows, as shown in Table II.
  • The impacts of JAD speed $v^{*}$ and wave width $\Delta_{\mathrm{w}}$ on $J_{\mathrm{dis}}$ and $J_{\mathrm{dur}}$ are positive, i.e., higher $v^{*}$ and larger $\Delta_{\mathrm{w}}$ lead to a longer JAD process, while fixing all other parameters.
  • The impacts of inflow traffic speed $v_{\mathrm{t}}$, wave speed $w$, and in-jam traffic speed $v_{\mathrm{w}}$ on $J_{\mathrm{dis}}$ and $J_{\mathrm{dur}}$ are negative, indicating that a lower $v_{\mathrm{t}}$, a faster propagating wave (smaller $|w|$), and lower $v_{\mathrm{w}}$ require a longer JAD process.
@@
  • Considering the moderate values of the other parameters, the relative influence of each parameter on $J_{\mathrm{dis}}$ and $J_{\mathrm{dur}}$, from strongest to weakest, is:
@@
1 JAD speed ($v^{*}$) ++ [2, 13] [50, 160]
2 Wave width ($\Delta_{\mathrm{w}}$) ++ [1, 8] [30, 140]
3 Wave speed ($w$) - [3, 9] [50, 150]
4 Inflow speed ($v_{\mathrm{t}}$) - [4, 7.5] [60, 130]
5 In-wave speed ($v_{\mathrm{w}}$) - [4, 5.4] [60, 90]
@@ V-A Simulation Scenario
It is known that speed variance between lanes is usually small, as lane-changing behavior tends to smooth out inter-lane differences [11, 25, 26]. Therefore, single-lane traffic or the average across all lanes is representative of multi-lane traffic. Without loss of generality, we simulate traffic on a single-lane straight freeway. SUMO is employed to simplify reproducibility. The road length is 8 km, with a speed limit of 90 km/h. The simulation runs for 1600 s with a time step of 1 s. Vehicles enter the network at the freeway entrance every 1.5 s [^4] Such a persistently high demand is effective in ensuring the continuous growth of a stop-and-go wave. We also experimented with more stochastic demand patterns, but failed to produce a self-sustaining wave. with an initial speed of 54 km/h, and they soon accelerate to their maximum speed. The Krauss car-following model embedded in SUMO is employed, and the main parameters are summarized in Table III.
@@
Figure 10: SUMO simulatied self-sustaining stop-and-go wave.
@@ V-B1 JAD Speed ($v^{*}$)
We reduce the speed of the first vehicle to 30, 40, 50, and 60 km/h, respectively, and simulate each scenario to observe which speeds do not trigger secondary waves. As shown in Figure 11, speeds of 30 and 40 km/h result in traffic breakdown, and are therefore unsuitable as JAD speeds. In contrast, speeds of 50 km/h or higher can serve as reliable JAD speeds. However, as discussed in Section IV-A, increasing the JAD speed also increases the time and space required to execute the JAD task, implying a longer operational period.
@@
V-B2 Wave temporal width ($\Delta_{\mathrm{w}}$)
@@
Typically, a stationary detector monitors the traffic and measures the speed of passing vehicles in real time. Therefore, we assume that a stop-and-go wave has formed when the observed speed falls below a certain threshold and continues to remain low for a period of time. This approach is feasible because the target of this study is an isolated stop-and-go wave. In contrast, for periodic stop-and-go waves generated by fixed bottlenecks, this method may not be suitable, and the proposed JAD strategy is not designed for such traffic waves.
In this scenario, we set the threshold to 36 km/h and the holding period to 30 s. As shown in Figure 12, the resulting time-space points $F$ and $E$ effectively capture the stop-and-go wave. It is also noted that, due to the asymmetric behavior of vehicles traversing a stop-and-go wave [4, 17, 23], the recovery process from a slowly moving queue is longer than the deceleration process. As will be demonstrated later in Section V-C with simulation results, a slight overestimation of the wave speed, while probably sacrificing some efficiency, not only does not compromise the goal of halting the propagation of the stop-and-go wave, but may also improve the robustness of the JAD strategy. If one truly wants to avoid generating such an artificial gap in traffic flow, a more advanced strategy such as the two-step strategy proposed in [8] can be adopted. However, it should be noted that more real-time traffic information is typically required to support such fine-grained management.
@@
V-B3 In-wave Speed ($\Delta_{\mathrm{w}}$)
From the practical perspective of a stationary roadside detector, the speed variation through a stop-and-go wave exhibits a gradual decrease followed by a gradual increase (also illustrated in Figure 12). As a result, the stationary roadside detector may be unable to accurately capture the in-wave speed.
Here, we take the average speed between $t_{F}$ and $t_{E}$ as the in-wave speed $\Delta_{\mathrm{w}}$. If the minimum value were used instead, $\Delta_{\mathrm{w}}$ could be underestimated.
V-B4 Wave Speed ($w$)
@@
Here, from the simulation, we set the wave speed to $w$ = -15 km/h.
@@ V-B7 Opportunity of Dispatching a JAD Vehicle
In the police-car swerving scenario, the police officer is expected to identify an appropriate gap for merging. If merging from an on-ramp, the overall process would be as natural as typical daily freeway merging behavior. With the siren activated, the required gap may be even smaller than that under normal conditions.
@@ V-C Simulation Results
To moderate multi-lane traffic, swerving behavior is required. Since such behavior has already been observed in the real world (Figure 1), it is unnecessary to simulate or verify its feasibility. Therefore, we focus on evaluating the stop-and-go mitigation effect of the JAD strategy rather than the specific swerving maneuvers of a JAD vehicle.
@@
The results could be further improved by fine-tuning the parameters. For example, a more accurate estimation of the wave width could enhance efficiency by enabling a better prediction of the ending point $C$ of the JAD strategy. However, this refinement may not be easy to implement, because, in practice, traffic conditions cannot be perfectly replicated, and obtaining an exact value is challenging.
Note that it is not necessary to compare travel time or emissions with the baseline in Figure 10, since once the stop-and-go wave is suppressed, the following vehicles naturally experience shorter travel times and lower emissions when operating under smoother traffic conditions without falling into a stop-and-go process.
@@
Figure 15 presents two typical cases in which the JAD vehicle fails to suppress the stop-and-go wave, as a secondary wave is triggered. In Figure 15(a), the JAD speed is set to $v^{*}$ = 35 km/h. As indicated by the stability test in Figure 11, such a low speed is insufficient to prevent traffic breakdown. In Figure 15(b), we underestimate the wave width intentionally, and the JAD vehicle fails to avoid being captured by the stop-and-go wave. As a result, the stop-and-go wave penetrates the JAD trajectory and continues to propagate upstream.